# Energy Transduction in Complex Networks with Multiple Resources: The Chemistry Paradigm

### v1

@@ Introduction
The first part of this work addresses this gap by outlining the key requirements for analyzing transduction and establishing a systematic procedure for selecting the correct set of processes. This procedure has two key conceptual highlights: First, it generalizes the framework for defining transduction in thermal engines connected to multiple heat baths [17]. As in that context, a reference equilibrium environment is specified, which defines the chemical exergy—that is, the maximum work extractable from out-of-equilibrium chemostats relative to that reference equilibrium. By adopting this approach, we adhere to a physically meaningful definition of efficiency, expressed as the ratio of output to input exergy. Second, we analyze transduction in terms of elementary processes—those that cannot be further decomposed into subprocesses—a new concept introduced in this work. These processes represent the smallest fundamental units of free energy that can be produced or consumed in the chemostats. By basing our analysis on elementary processes, we ensure that all transduction occurring in the CRN is captured. Once the set of processes is selected, we prescribe a method for classifying them as input or output processes, enabling a well-defined measure of transduction efficiency. Unlike previous studies, our framework allows for the exclusion of ‘unusable’ outputs, accounting for potential downstream losses that are not captured within the CRN. Additionally, we emphasize the inherently relative nature of transduction efficiency, as it may depend on the choice of the reference equilibrium, the definition of ‘useful’ outputs, and the scope of the CRN analyzed.
@@ I.1 Setup
We consider open and well-stirred chemical reaction networks (CRNs) under stationary conditions, consisting of a set of reversible reactions $\rho$, internal species $X$, and external species $Y$. The external species are chemostatted, meaning they are continuously exchanged with the environment, which ensures that their concentrations—and thus their chemical potentials—remain constant over time. Figures 1a and 1c illustrate the two representative open CRNs analyzed in this paper. The first depicts a network combining glycolysis and gluconeogenesis (GG network) [16], while the second represents a coarse-grained yet thermodynamically complete model of yeast metabolism capturing both respiration and fermentation (RF network) [21]. As detailed in Sects. I.3 and I.4, this simplified representation retains all necessary information for our analysis. In both cases, the $Y$ species are highlighted. The state of the system is defined by the vector of concentrations
@@
where $\boldsymbol{x}=(\dots,[x_{i}],\dots)$ denotes the concentrations of the internal species $X$ and $\boldsymbol{y}=(\dots,[y_{i}],\dots)$ denotes the concentrations of the external species $Y$. The evolution of these concentrations is governed by the stoichiometric matrix $\mathbb{S}$ [25], which defines the topology of the CRN by encoding how reactions interconvert species. By separating the stoichiometric matrix with respect to internal and external species,
@@ I.3 Atomic Composition Matrix
In what follows, we focus exclusively on effective reactions among external species $Y$. Accordingly, only $\mathbb{A}^{Y}$, which encodes the atomic composition of external species, is relevant. Therefore, a coarse-graining of internal species or reactions is admissible, provided the resulting network gives rise to effective reactions among external species that preserve mass and charge—as in the RF network in Fig. 1c. In addition, the coarse-graining must maintain thermodynamic consistency, as discussed in the next section.
@@ II.1 Definition
Elementary processes constitute a special class of processes representing the most basic interconversions among the $Y$ species. These are defined as processes that cannot be further decomposed into subprocesses (see Appendix A for more details). For example, $\boldsymbol{p}_{\epsilon_{1}}^{\text{GG}}$ is elementary, while $\boldsymbol{p}_{\epsilon_{1}}^{\text{RF}}$ is not. The reason is that the latter carries out two subprocesses: the conversion of glucose to ethanol, $\text{Glc}\rightarrow 2\text{Eth}+2\text{CO}_{2}$, and ATP synthesis, $2\text{ADP}+2\text{P}_{i}+2\text{H}^{+}\rightarrow 2\text{ATP}+2\text{H}_{2}\text{O}$, both of which are elementary.
@@ II.2 Complete and Independent Sets of Processes
This section aims to identify a set of processes that provides an unambiguous description of matter flow through the CRN, that is, of the intake and release of $Y$ species quantified by the exchange flux vector $\boldsymbol{I}^{Y}$ in Eq. (5b). Establishing such a set will be essential for the later discussion of transduction in Sect. III.1. To serve this purpose, the set of processes must meet two fundamental requirements: completeness and linear independence. We call a set $\{\boldsymbol{p}\}$ complete if linear combinations of its elements can represent any process realizable by the CRN (i.e., any process in $\Pi_{\text{CRN}}$ defined in Eq. (23)). Such completeness enables the decomposition of $\boldsymbol{I}^{Y}$:
@@
where $\mathcal{I}_{\boldsymbol{p}}$ denotes the steady state rate at which process $\boldsymbol{p}$ occurs: $\mathcal{I}_{\boldsymbol{p}}>0$ when the process proceeds forward and $\mathcal{I}_{\boldsymbol{p}}<0$ when it proceeds backward. If the set $\{\boldsymbol{p}\}$ is both complete and linearly independent, this decomposition becomes unique, enabling a corresponding unique decomposition of the entropy production in the chemostats:
@@
Here, $\boldsymbol{\mu}_{Y}$ represents the vector of chemical potentials of the $Y$ species, which specifies the operating conditions, and $\Delta_{\boldsymbol{p}}G=\boldsymbol{\mu}_{Y}\cdot\boldsymbol{p}$ is the Gibbs free energy change caused by process $\boldsymbol{p}$. This Gibbs free energy change is equivalent to the sum of the products’ chemical potentials minus the sum of the reactants’ chemical potentials. Examples of complete and independent sets of processes for the GG and RF networks will be presented later in Eqs. (34) and (35), together with the corresponding decompositions of $-\boldsymbol{I}^{Y}$ and $\dot{\Sigma}$ in Eqs. (37) and (38).
@@ III.4 Definition of Efficiency
Once a proper set of processes $\mathcal{T}$ has been identified, we can define the transduction efficiency $\eta$. Here, the key conceptual step is determining which processes qualify as input and which as output. Notably, this classification, $\mathcal{T}=\mathcal{T}_{\text{in}}\cup\mathcal{T}_{\text{out}}$, is not solely dictated by the network topology and operating conditions but also by the specific CRN stationary state. In particular, the reaction fluxes within the network determine whether a process proceeds thermodynamically downhill (input) or uphill (output). This classification naturally emerges from entropy production, where positive contributions correspond to input processes and negative contributions to output processes. We first illustrate the procedure for the two CRN examples before providing the general definition in Eq. (53). From Eq. (38a), the stationary entropy production of the GG network decomposed in terms of $\mathcal{T}^{\text{GG}}$ reads
@@
Using this last example as an illustration, we highlight an important feature of the framework: the flexibility to selectively include only a portion of the output processes when defining efficiency. This is relevant in cases where certain output processes are ‘unusable,’ as the free energy they store is ultimately dissipated through reactions occurring outside the CRN, making it effectively irrecoverable. For example, consider a scenario in which a yeast (RF network) ferments glucose into ethanol, and bacteria in the surrounding environment consume that ethanol as part of subsequent stages in the ecological chain. From the yeast’s perspective, the free energy stored in ethanol is effectively lost, as the yeast does not reuse it. As a result, the efficiency with which the yeast captures the glucose’s free energy relies solely on the ATP produced during fermentation. This efficiency is therefore smaller and equal to
@@
While focusing on a subset of output processes ensures that the resulting efficiency remains below 1, the same cannot be said for input processes. Input processes represent free energy costs in the system, and neglecting any of these costs could result in an unphysical efficiency bigger than 1.
@@
    Depending on the system under analysis, restrict the set of output processes, considering only the ‘useful’ ones $\mathcal{T}_{\text{out}}^{*}\subset\mathcal{T}_{\text{out}}$:
@@ V.2 Applications
The gears conformal to these flux directions are $\boldsymbol{\psi}_{\alpha},\,\boldsymbol{\psi}_{\beta},\,\boldsymbol{\psi}_{\gamma}$. Experimental evidence shows that shifting the operating conditions toward the lower left in Fig. 3e—i.e., increasing glucose concentration—drive yeast to upregulate fermentative metabolism (gear $\boldsymbol{\psi}_{\gamma}$), a phenomenon known as the Crabtree effect [21, 36]. Second, within the physiological range, fermentation—mediated by gears $\boldsymbol{\psi}_{\gamma}$ and $\boldsymbol{\psi}_{\delta}$—is more efficient than respiration, which relies on gears $\boldsymbol{\psi}_{\alpha}$ and $\boldsymbol{\psi}_{\beta}$ (when ethanol is considered as a ‘useful’ output). Notably, fermentation achieves an efficiency exceeding 90%, aligning with experimental observations [37]. This finding underscores the importance of distinguishing thermodynamic efficiency from stoichiometric yield, a commonly used metric quantifying the number of ATP molecules produced per glucose molecule consumed. Although the stoichiometric yield of fermentation is significantly lower than that of respiration—since fermentation captures only a small fraction of glucose’s energy as ATP—its thermodynamic efficiency is higher.

### v2

@@ Introduction
The first part of this work addresses this gap by outlining the key requirements for analyzing transduction and establishing a systematic procedure for selecting the correct set of processes. This procedure has two key conceptual highlights: First, it generalizes the framework for defining transduction in thermal engines connected to multiple heat baths [17]. As in that context, a reference equilibrium environment is specified, which defines the chemical exergy—that is, the maximum work extractable from out-of-equilibrium chemostats relative to that reference equilibrium. By adopting this approach, we adhere to a physically meaningful definition of efficiency, expressed as the ratio of output to input exergy. Second, we analyze transduction in terms of elementary processes—those that cannot be further decomposed into subprocesses—a new concept introduced in this work. These processes represent the smallest fundamental units of free energy that can be produced or consumed in the chemostats. They are defined exclusively by the atomic composition of chemical species, using the atomic composition matrix as the primary analytical tool. By basing our analysis on elementary processes, we ensure that all transduction occurring in the CRN is captured. Once the set of processes is selected, we prescribe a method for classifying them as input or output processes, enabling a well-defined measure of transduction efficiency. Unlike previous studies, our framework allows for the exclusion of ‘unusable’ outputs, accounting for potential downstream losses that are not captured within the CRN. Additionally, we emphasize the inherently relative nature of transduction efficiency, as it may depend on the choice of the reference equilibrium, the definition of ‘useful’ outputs, and the scope of the CRN analyzed.
@@ I.1 Setup
We consider open and well-stirred chemical reaction networks (CRNs) under stationary conditions, consisting of a set of reversible reactions $\rho$, internal species $X$, and external species $Y$. All reactions are mass- and charge-preserving conversions among species. The external species are chemostatted, meaning they are continuously exchanged with the environment, which ensures that their concentrations—and thus their chemical potentials—remain constant over time. Figures 1a and 1c illustrate the two representative open CRNs analyzed in this paper. The first depicts a network combining glycolysis and gluconeogenesis (GG network) [16], while the second represents a coarse-grained yet thermodynamically complete model of yeast metabolism capturing both respiration and fermentation (RF network) [21]. As detailed in Sects. I.3 and I.4, this simplified representation retains all necessary information for our analysis. In both cases, the $Y$ species are highlighted. The state of the system is defined by the vector of concentrations
@@
where $\boldsymbol{x}=(\dots,[x_{i}],\dots)$ denotes the concentrations of the internal species $X$ and $\boldsymbol{y}=(\dots,[y_{i}],\dots)$ denotes the concentrations of the external species $Y$. The evolution of these concentrations is governed by the stoichiometric matrix $\mathbb{S}$ [25], which encodes the topology of the CRN by specifying how reactions interconvert species. Defining this matrix requires establishing the forward direction for each reaction. By separating the stoichiometric matrix with respect to internal and external species,
@@ I.3 Atomic Composition Matrix
In what follows, we focus exclusively on effective reactions among external species $Y$ since, under stationary conditions, they are the only ones being produced or consumed. Accordingly, only $\mathbb{A}^{Y}$, which encodes the atomic composition of external species, is relevant. Therefore, a coarse-graining of internal species or reactions is admissible, provided the resulting network gives rise to effective reactions among external species that preserve mass and charge—as in the RF network in Fig. 1c. In addition, the coarse-graining must maintain thermodynamic consistency, as discussed in the next section.
@@ II.1 Definition
Elementary processes constitute a special class of processes representing the most basic interconversions among the $Y$ species. These are defined as processes that cannot be further decomposed into subprocesses (see Appendix A for more details including how to compute them). For example, $\boldsymbol{p}_{\epsilon_{1}}^{\text{GG}}$ is elementary, while $\boldsymbol{p}_{\epsilon_{1}}^{\text{RF}}$ is not. The reason is that the latter carries out two subprocesses: the conversion of glucose to ethanol, $\text{Glc}\rightarrow 2\text{Eth}+2\text{CO}_{2}$, and ATP synthesis, $2\text{ADP}+2\text{P}_{i}+2\text{H}^{+}\rightarrow 2\text{ATP}+2\text{H}_{2}\text{O}$, both of which are elementary.
@@ II.2 Complete and Independent Sets of Processes
This section aims to identify a set of processes that provides a univocal description of matter flow through the CRN, that is, of the intake and release of $Y$ species quantified by the exchange flux vector $\boldsymbol{I}^{Y}$ in Eq. (5b). Establishing such a set will be essential for the later discussion of transduction in Sect. III.1. To serve this purpose, the set of processes must meet two fundamental requirements: completeness and linear independence. We call a set $\{\boldsymbol{p}\}$ complete if linear combinations of its elements can represent any process realizable by the CRN (i.e., any process in $\Pi_{\text{CRN}}$ defined in Eq. (23)). Such completeness enables the decomposition of $\boldsymbol{I}^{Y}$:
@@
where $\mathcal{I}_{\boldsymbol{p}}$ denotes the steady state rate at which process $\boldsymbol{p}$ occurs: $\mathcal{I}_{\boldsymbol{p}}>0$ when the process proceeds forward and $\mathcal{I}_{\boldsymbol{p}}<0$ when it proceeds backward. If the set $\{\boldsymbol{p}\}$ is both complete and linearly independent, this decomposition becomes unique. Substituting Eq. (24) into the entropy production expression for the chemostats then yields a corresponding unique decomposition of the entropy production:
@@
Here, $\boldsymbol{\mu}_{Y}$ represents the vector of chemical potentials of the $Y$ species, which specifies the operating conditions, and $\Delta_{\boldsymbol{p}}G=\boldsymbol{\mu}_{Y}\cdot\boldsymbol{p}$ is the Gibbs free energy change caused by process $\boldsymbol{p}$. This Gibbs free energy change is equivalent to the sum of the products’ chemical potentials minus the sum of the reactants’ chemical potentials. For any given CRN, it is always possible to identify a complete and independent set of processes. Examples for the GG and RF networks will be presented later in Eqs. (34) and (35), together with the corresponding decompositions of $-\boldsymbol{I}^{Y}$ and $\dot{\Sigma}$ in Eqs. (37) and (38).
@@ III.4 Definition of Efficiency
Once a proper set of processes $\mathcal{T}$ has been identified, we can define the transduction efficiency $\eta$. Here, the key conceptual step is determining which processes qualify as input and which as output. Notably, this classification, $\mathcal{T}=\mathcal{T}_{\text{in}}\cup\mathcal{T}_{\text{out}}$, is not solely dictated by the network topology and operating conditions but also by the specific CRN stationary state. In particular, the reaction fluxes within the network determine whether a process proceeds thermodynamically downhill (input) or uphill (output). This classification naturally emerges from entropy production, where positive contributions correspond to input processes and negative contributions to output processes. We first illustrate the procedure for the two CRN examples before providing the general definition in Eq. (53).
Examples.
From Eq. (38a), the stationary entropy production of the GG network decomposed in terms of $\mathcal{T}^{\text{GG}}$ reads
@@
Using this last example as an illustration, we highlight an important feature of the framework: the flexibility to selectively include only a portion of the output processes when defining efficiency. This is particularly valuable when certain outputs, though thermodynamically significant, are not utilized by the system—their stored free energy is ultimately dissipated through external reactions, making it effectively irrecoverable. For example, consider a scenario in which a yeast (RF network) ferments glucose into ethanol, and bacteria in the surrounding environment consume that ethanol as part of subsequent stages in the ecological chain. From the yeast’s perspective, the free energy stored in ethanol is effectively lost, as the yeast does not reuse it. As a result, the efficiency with which the yeast captures the glucose’s free energy relies solely on the ATP produced during fermentation. This efficiency is therefore smaller and equal to
@@
To formalize this approach, depending on the specific system, one can partition the full set of output processes into two disjoint subsets: target output processes (or ‘useful’ processes)—those whose stored free energy can be harnessed by the system at a later stage—and non-recoverable output processes, whose stored free energy is effectively lost. This distinction allows us to define the target efficiency, denoted by $\eta^{*}$, which quantifies the system’s effectiveness at converting input free energy solely into its designated target outputs. We formalize it later in Eq. (54). We note that while focusing on a subset of output processes ensures that the resulting target efficiency remains below 1, the same cannot be said for input processes. Input processes represent free energy costs in the system, and neglecting any of these costs could result in an unphysical efficiency bigger than 1.
General Procedure
@@
    Depending on the system under analysis, restrict the set of output processes only to the useful ones, $\mathcal{T}_{\text{out}}^{*}\subset\mathcal{T}_{\text{out}}$:
@@ V.2 Applications
The gears conformal to these flux directions are $\boldsymbol{\psi}_{\alpha},\,\boldsymbol{\psi}_{\beta},\,\boldsymbol{\psi}_{\gamma}$. Experimental evidence shows that shifting the operating conditions toward the lower left in Fig. 3e—i.e., increasing glucose concentration—drive yeast to upregulate fermentative metabolism (gear $\boldsymbol{\psi}_{\gamma}$), a phenomenon known as the Crabtree effect [21, 36]. Second, within the physiological range, fermentation—mediated by gears $\boldsymbol{\psi}_{\gamma}$ and $\boldsymbol{\psi}_{\delta}$—is more efficient than respiration, which relies on gears $\boldsymbol{\psi}_{\alpha}$ and $\boldsymbol{\psi}_{\beta}$ (when ethanol is considered as a useful output). Notably, fermentation achieves an efficiency exceeding 90%, aligning with experimental observations [37]. This finding underscores the importance of distinguishing thermodynamic efficiency from stoichiometric yield, a commonly used metric quantifying the number of ATP molecules produced per glucose molecule consumed. Although the stoichiometric yield of fermentation is significantly lower than that of respiration—since fermentation captures only a small fraction of glucose’s energy as ATP—its thermodynamic efficiency is higher.
@@ Acknowledgements
VII Author contributions
MB and ME contributed equally to all aspects of this work, including the conceptualization, design, analysis, and writing of the manuscript.
VIII Conflict of interest
The authors declare that they have no conflicts to disclose.
IX Data Availability
The data that support the findings of this study are available within the article.